Part A Calculate the force of Earths gravity on a spacecraft 2.00 Earth radii above the Earths surface if its mass is 1550
Answer #1

Given mass of spacecraft, m = 1550kg

Altitude of spacecraft, h = 2RE , where R_{E} is radius of earth.

So Distance of spacecraft from center of earth, D=RE +h=RE +2RE = 3RE

D = 3R

------------------

Mass of Earth, ME = 5.972 * 104 kg

Radius of Earth,RE= 6.371 * 100m

--------------

Force of Earth's gravity on spacecraft, FG = Glem 02

F_{G}=\frac{GM_{E}m}{(3R_{E})^{2}}

GMem FG = 9RE

F_{G}=\frac{6.674*10^{-11}*5.972*10^{24}kg*1550kg}{9*(6.371*10^{6}m)^{2}}

F_{G}=\frac{6.674*10^{-11}*5.972*10^{24}*1550}{9*6.371^{2}*10^{12}}

F_{G}=\frac{6.674*5.972*1550*10^{(-11+24-12)}}{9*6.371^{2}}

FG = - 6.674 * 5.972 * 1550 * 101 9* 6.3712

ANSWER :{\color{Red} F_{G}=1671.142N}

===========================